Mr. Chen needs to raise capital for his business. He decides his capital need is $1,000,000. He wants to raise 40% in debt and 60% in equity. He will offer five year 5% bonds and common stock at $100 per share. Mr. Bedford buys all the bonds and all the stock on January 1. 1. Record the transaction on the books of Mr. Chen and and on Mr. Bedfords books. 2. Record the interest payment made on 7/1 on the books of Mr. Chen and Mr. Bedford. If one year later bods rated at a similar degree of risk to the Chen bonds are paying: A) 4% B) 9% Calculate what you will pay for the Chen Bonds under assumption A and B. Record the purchase of the bonds under each assumption

steo by step please
Answer #1
  1. Accounting Entries for Bond Issue

                        In the books of Chen

                       

Cash/ Bank A/c ……………………Dr                           4,00,000

To Bond Payable A/c                                                                                   4,00,000

In the books of Mr Bedford

Investment in Bond A/c ………….Dr                        4,00,000

To Cash/ Bank                                                                                             4,00,000

2.Interest Payment

            In the books of Chen

Interest Expense A/c ………Dr                                 10000

To Cash / Bank                                                                                            10000

In the books of Bedford

Cash / Bank A/c ………….Dr                                      10000

To Interest Income A/c                                                                            10000

Note: Since the interest payment is on July 1 it assuming that it will be pay half yearly

So Interest = 40000*5%*6/12=10000. Since the question asked to record only the interest payment made on 7/1, the next half year interest is not recorded here.

Answer for the Second Part of the Question

After one year bond rated at similar degree of risk to the Chen bond are paying

A) 4%

Then the Payment for the Chen bond will be as follows

For this purpose we have to compute the Present value of the Interest for the remaining years and Present value of bond principal amount

The fact of the question is as follows

Remaining years for Mature the bond = 5year – 1 year = 4 year

Coupon rate for computing present value is market rate , that is 4%

Since interest payment is semi- annually we should take tthe coupon rate 2% for 8 years.

  

Value of the Bond = Present Value of Interests + present Value of principal amount of bond

Present value of Interest = Interest* Annuity factor of 2 % for 8 Years

=10000*7.3255

=73255

Present value of principal amount of Bond= principal Amount repaying*PV factor for the last year ( In the case forth year after one year)

That is PV factor of 2 % in year 8

=400000*.8535 =341400

Hence We will pay for the Chen Bond under Assumption A

73255+341400 = 4,14,655

Assumption B- 9%

As the same step followed in Assumption A

Value of the bond = Present value of interest + Present Value of Principal

For calculating Present value / Annuity Factor the rate is 4.5% that is ( 9/2) for 8 years

Present value of Interest =10000*6.5958 =65959

Present value of Principal amount =400000*0.7032 =281280

Value of Bond =65959+281280

Hence we will pay 3,47,239 for Chen bond under assumption B

Entry For Purchase of Bond

Assumption A

In this case the bond is under priced since the value of the bond is 4,14,655 is more than the issue price, so it issuing at discount

Investment in Bond A/c …………………Dr4,14,655

To Cash / Bank4,00,00

To Discount on Bond14,655

Assumption B

In this case the bond is overpriced since the value of the bond is 3,47,239 is less than the issue price, so it is issuing at premium

Investment in Bond A/c ……………Dr 347239

Premium on Bond A/c ………………Dr 52761

To Cash/ Bank 4,00,000